If the ratio of y + 5x to x + 2y is 2/3, what is the ratio of y to x?
Answer by stanbon(48575)

If the ratio of
(y + 5x) : (x + 2y) is 2/3,
what is the ratio of y : x?
================================
Equation:
(y+5x/(x+2y) = 2/3
3y+15x = 2x+4y
y = 13x
y/x = 13/1
--------------------
y:x as 13:1
==========================
Cheers,
Stan H.